OpenStudy (anonymous):

Solve ^3sqrt(3x-11)= ^3sqrt(x+12) Is the answer 23/3 ?

OpenStudy (anonymous):

what have you tried ? :)

OpenStudy (anonymous):

what is this ^3?

OpenStudy (anonymous):



OpenStudy (anonymous):

is it that?

OpenStudy (anonymous):

\[\huge \sqrt[3]{3x-11}=\sqrt[3]{x+12}\]

OpenStudy (anonymous):

Yeah thats it

OpenStudy (anonymous):

so then it isnt sqrdrt its like triple root lol

OpenStudy (anonymous):

Just power both sides with 3 and you'll get 3x-11=x+12 --> 2x=23 --> x=23/2

OpenStudy (anonymous):

Almost right! :D

OpenStudy (anonymous):

... I didn't want to provide direct answer, let him solve it, let's see where he's stuck. :)